OpenStudy (anonymous):

If 4m-7n=10 and 2m+2n =4 what is the value of 2m-9n?

OpenStudy (amistre64):

they want you to find the common solution tothe first 2 equations and use those numbers in the last one

OpenStudy (a_clan):

or simply subtract second from first

OpenStudy (amistre64):

id multiply 2nd by -2 then add them together

OpenStudy (anonymous):

I dont really understand how to do that though... I've tried a lot of numbers for finding the solution to the first one, but it never worked out for the second one, or vice versa

OpenStudy (amistre64):

trial and error is a shot in the dark; you have to use numerical methods such as substitution or elimination

OpenStudy (a_clan):

subtract second equation from first equation

OpenStudy (amistre64):

4m-7n=10 2m+2n =4 <-- multiply this one by -2 4m-7n=10 -4m-4n =-8 ; now add them together ----------- 0 -11n = 2 , now solve for n

OpenStudy (anonymous):

oh... thank you

OpenStudy (amistre64):

once you know "n"; solving for m is easy :)